### v2

@@ 6 Related Work
Inspired by the success of diffusion models in continuous image domains Ho et al. (2020); Rombach et al. (2022), recent work has extended diffusion to discrete text generation. Early approaches operating in continuous embedding spaces Li et al. (2022); Gong et al. (2022); Han et al. (2022) suffered from optimization and discretization issues. In contrast, masked diffusion models Lou et al. (2023); Sahoo et al. (2024); Shi et al. (2024); Ou et al. (2024), which define the diffusion process directly in the discrete token space via random masking, have become the dominant paradigm. Notably, recent large-scale models such as LLaDA Nie et al. (2025) and Dream Ye et al. (2025) demonstrate scalability and performance competitive with autoregressive (AR) models. The emergence of such large-scale diffusion language models has spurred a growing body of follow-up studies exploring their distinctive modeling and inference properties. Among various advantages discussed in the literature, two aspects have received particular attention. First, dLLMs naturally support parallel decoding, enabling significant inference acceleration Wu et al. (2025b); Ben-Hamu et al. (2025); Labs et al. (2025); DeepMind (2025); Song et al. (2025). Second, their non-autoregressive formulation allows for arbitrary-order token generation, which has been hypothesized to benefit complex reasoning by relaxing strict left-to-right constraints Ye et al. (2024); Kim et al. (2025).
@@
While early studies validated the utility of arbitrary order generation in constrained tasks like Sudoku and Zebra Puzzles Ye et al. (2024); Nie et al. (2025); Kim et al. (2025), recent research has begun to investigate its value in general reasoning domains. One line of research, exemplified by Dream-Coder Xie et al. (2025) and DiffuCoder Gong et al. (2025), shows that diffusion models can naturally exhibit non-standard, human-like decoding behaviors (e.g., sketch-first reasoning) without any explicit supervision on generation order. DiffuCoder Gong et al. (2025) further quantifies this phenomenon by demonstrating that higher sampling temperatures reduce AR-ness, with increased order randomness correlating with improved output diversity. Another line of research, exemplified by P2 Peng et al. (2025) and LLaDOU Huang et al. (2025b), explicitly optimizes the decoding strategy for better generation order, thus better performance. Despite these advances, it remains unclear whether the observed gains primarily arise from better exploitation of existing solution patterns encoded in the data and model, or whether order arbitrariness itself enables qualitatively new reasoning strategies that are unattainable under a purely autoregressive decoding regime.
@@
Reinforcement learning for dLLMs faces structural optimization hurdles distinct from the autoregressive paradigm, primarily stemming from the combinatorial explosion of denoising trajectories. While early attempts Zhao et al. (2025); Yang et al. (2025); Gong et al. (2025) sought to adapt token-level formulations directly, they were fundamentally limited by ill-defined state transitions, necessitating reliance on unstable mean-field approximations. Consequently, the field has shifted toward sequence-level perspectives Zhu et al. (2025); Wang et al. (2025a); Rojas et al. (2025); Ou et al. (2025), employing various surrogates to approximate the intractable marginal likelihood. Yet, a critical off-policy misalignment persists across these methods: the heuristic-guided sampling required for efficient exploration diverges from the underlying diffusion prior, rendering gradients biased without principled correction Schulman et al. (2015). Two notable exceptions partially address these issues. LLaDOU Huang et al. (2025b) explicitly models token position selection via an auxiliary policy, enabling direct estimation of trajectory likelihoods, while TraceRL Wang et al. (2025d) aligns optimization with inference traces through a shrinkage-based step-wise MDP formulation. Nevertheless, these approaches remain committed to preserving the full arbitrary-order generation mechanism, implicitly treating its structural complexity as indispensable, whereas we question whether effective RL training can be achieved more simply by reexamining the necessity of arbitrary order itself.

### v3

@@ 6 Related Work
Motivated by the success of diffusion models in continuous domains Ho et al. (2020); Rombach et al. (2022), recent work has extended diffusion to discrete text generation. Early embedding-space approaches Li et al. (2022); Gong et al. (2022); Han et al. (2022) faced optimization and discretization challenges, leading to the adoption of masked diffusion models Lou et al. (2023); Sahoo et al. (2024); Shi et al. (2024); Ou et al. (2024), which operate directly in the discrete token space via random masking. Recent large-scale models like LLaDA Nie et al. (2025) and Dream Ye et al. (2025) achieve performance competitive with autoregressive (AR) models. Among various advantages of dLLMs discussed in the literature, two aspects have received particular attention. First, dLLMs naturally support parallel decoding, enabling significant inference acceleration Wu et al. (2025b); Ben-Hamu et al. (2025); Labs et al. (2025); DeepMind (2025); Song et al. (2025). Second, their non-autoregressive formulation allows for arbitrary-order token generation, which has been hypothesized to benefit complex reasoning by relaxing strict left-to-right constraints Ye et al. (2024); Kim et al. (2025).
@@
While early studies validated arbitrary order generation in constrained tasks Ye et al. (2024); Nie et al. (2025); Kim et al. (2025), recent research extends this to general reasoning. Some works demonstrate that dLLMs can naturally exhibit non-standard decoding behaviors (e.g., sketch-first) Xie et al. (2025); Gong et al. (2025), while others explicitly optimize decoding schedules to boost performance Peng et al. (2025); Huang et al. (2025b). Moreover, some studies Gong et al. (2025); Hong et al. (2025); Shen et al. (2026) also employ Pass@$k$ to demonstrate the reasoning potential of dLLMs. Recognizing this potential, these works predominantly focus on enhancing performance within the arbitrary-order paradigm, while leaving whether the benefit is uniquely tied to the order arbitrariness less-examined. Recently, Du et al. (2025) challenged the necessity of arbitrary order, arguing that forcing models to learn uniform permutations causes a significantly looser approximation of the underlying data distribution. While their analysis focuses on pre-training, we identify a parallel failure mode in reasoning and RL, revealing how the flexibility degrades exploration essential for RL training.
@@
Reinforcement learning for dLLMs faces structural optimization hurdles distinct from the autoregressive paradigm, primarily stemming from the combinatorial explosion of denoising trajectories. While early attempts Zhao et al. (2025); Yang et al. (2025); Gong et al. (2025) sought to adapt token-level formulations directly, they were fundamentally limited by ill-defined state transitions, necessitating reliance on unstable mean-field approximations. Consequently, the field has shifted toward sequence-level perspectives Zhu et al. (2025); Wang et al. (2025a); Rojas et al. (2025); Ou et al. (2025), employing various surrogates to approximate the intractable marginal likelihood. Yet, a critical off-policy misalignment persists across these methods: the heuristic-guided sampling required for efficient exploration diverges from the underlying diffusion prior, rendering gradients biased without principled correction Schulman et al. (2015). Two notable exceptions partially address these issues. LLaDOU Huang et al. (2025b) introduces an auxiliary policy to model token position selection, enabling direct trajectory likelihood estimation, while TraceRL Wang et al. (2025d) aligns optimization with inference traces via a shrinkage-based step-wise MDP. Nevertheless, both retain the full arbitrary-order mechanism, implicitly assuming its necessity, whereas we question whether effective RL training can be achieved more simply by reconsidering the need for arbitrary order itself.